The publishing company ReadIt publishes its magazine Survive Economics on the internet. In order to access it, readers have to purchase an annual subscription that is currently priced at P1 = 100 EUR. At this price, the company recorded Q1 = 600 subscriptions. The cost of providing the magazine online is independent of the number of subscriptions. . Recently, an analysis of the demand for Survive Economics lead to the trustworthy information that demand is linear and that the firm sells Q2 = 800 subscriptions if it sets the price at P2 = 50 EUR and that it sells Q3 = 200 subscriptions if it charges P3 = 200 EUR. The management knows that there is an ambiguous effect on subscription revenue if it increases or decreases the subscription price that is somehow related to the economic concept of a price elasticity of demand. In order to figure out up to what level it should increase or decrease its price or even leave it unchanged, it decided to employ the management consultancy Clever&Smart.
1
Expert's answer
2018-11-14T14:48:09-0500

P1 = 100 EUR, Q1 = 600 subscriptions.

P2 = 50 EUR, Q2 = 800 subscriptions,

P3 = 200 EUR, Q3 = 200 subscriptions.

Between P1 and P2 the price elasticity of demand is:

Ed = (800 - 600)/(50 - 100)*(50 + 100)/(800 + 600) = -3/7, so the demand is inelastic and the total revenue will decrease, if we decrease the price.

Between P1 and P3 the price elasticity of demand is:

Ed = (200 - 600)/(200 - 100)*(200 + 100)/(200 + 600) = 1.5, so the demand is elastic and the total revenue will increase, if we increase the price until point-price elasticity of demand Ed = 1.